OpenStudy (anonymous):

The table below shows the ticket rates for whale watching trips offered by Beluga Fun Tours. Age Ticket Price Under three years free Three to twelve years $34 Over twelve years $45 On a certain day, the company took 103 people on whale watching trips. There were 64 children aged twelve and under, of which some children were under three years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day? (64 – x)34 + (103 – 64) ∙ 45 = C (103 + x)34

OpenStudy (bahrom7893):

the top one

OpenStudy (bahrom7893):

(103-64)*45 + (64-x)*34 = C, yea that's right, the top one

OpenStudy (anonymous):

(64 – x)34 + (103 – 64) ∙ 45 = C (103 + x)34 + (103 – (64 + x)) ∙ 45 = C (103 - x)34 + (103 – 64) ∙ 45 = C (64 + x)34 + (103 – (64 + x)) ∙ 45 = C

OpenStudy (bahrom7893):

(64 – x)34 + (103 – 64) ∙ 45 = C